OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -4 and a vertex at (-7, -9)?

OpenStudy (anonymous):

Help

hero (hero):

If the roots of the quadratic are \(-10\) and \(-4\), then \(x + 10 = 0\) \(x + 4 = 0\) And \((x + 10)(x - 4) = 0\) Multiply the left side of the quadratic equation to get the form \(ax^2 + bx + c = 0\) Then use \(x = -\dfrac{b}{2a}\) to find the \(x\) coordinate of the vertex.

hero (hero):

Actually the quadratic equation should be \((x+10)(x + 4) = 0\)

OpenStudy (anonymous):

so it would be x^2+14x+40

hero (hero):

Yes, now verify that the vertex is (-7,-9)

OpenStudy (anonymous):

wait so x^2+14x+40 is my final answer

OpenStudy (anonymous):

?

hero (hero):

Only after you have verified the vertex.

OpenStudy (anonymous):

how do i do that

hero (hero):

Find the x coordinate using the formula I gave you above.

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

thts wat i got

hero (hero):

Good now, plug that into the quadratic expression, then simplify

OpenStudy (anonymous):

i got -9

hero (hero):

Very good

OpenStudy (anonymous):

so my final answer is : the quadratic function that is created with roots -10 and -4 and a vertex at (-7, -9) is x^2+14x+40

hero (hero):

Yes, and of course include all the steps you did to arrive at that conclusion.

OpenStudy (anonymous):

ok thank u so much for ur help

hero (hero):

you're welcome